QUESTION 25



Given the linear model -8x + 7y = 9, what value of y is predicted when x=-4? Round to the nearest hundredth.

QUESTION 26

:

The following linear model is created: y = 5x+15. What is the predicted value of the dependent variable when the independent variable has a value of 1.4?

QUESTION 27


The slope, to the nearest hundredth, of the line 4x + (-5)y = -3 is

QUESTION 28



Given the linear model y = 2x + 12 , find the residual when x=9 and the actually data point is found to be (9, 36). Round to the nearest hundredth.

QUESTION 29



Given the linear model -3x - 11y = -(-8), what value of y is predicted when x=4? Round to the nearest hundredth.

QUESTION 30



Given the linear model -8x - 10y = -(6), the "Fit" when x=-9 would be...? Round to the nearest hundredth.

QUESTION 31


Given the linear model -4x - 8y + -7 = 0, find the residual when x=-9 and the actually data point is found to be (-9, 8). Round to the nearest hundredth.
Answer #1

Note : As per Chegg policy, if a post contains more than one distinct questions, we should solve the first question or the question number suggested by the student.

Yet solving the question 25, 26, 27, 28

Q25)

-8x + 7y = 9

x = -4, Substituting value of x in the linear model

-8(-4) + 7y = 9

y = (9 - 32) / 7

y = -3.29

Q26)

y = 5x+15

x is the independent variable

y is the dependent variable

Given x = 1.4

Substituting value of x in the linear model

y = 5(1.4) + 15

y = 22

Predicted value of the dependent variable = 22

27)

4x + (-5)y = -3

Slope is the coefficient of the independent variable x

4x + (-5)y = -3

5y = 4x + 3

y = 0.8x + 0.6

Slope = 0.80

28)

y = 2x + 12

We find predicted value of y when x = 9

y = 2(9) + 12

y = 30

Actual data point is (9, 36)

Hence actual y = 36

Residual = Actual - Predicted

Residual = 36 - 30

Residual = 6      when x = 9